2007-03-08 ALMA Calibration Group Telecon Agenda and Minutes
Table of Contents:
ALMA Calibration Group Telecon |
- Date: March 8, 2007 (Thursday)
- Time: 15:00 UT
- Duration: 1.0 hr
- USA Toll Free Number: 1-877-919-7148
- International: +1-203-566-1039
- Passcode: 5 1 0 4 6 8
- Leader: JeffMangum
Action Items from Last Telecon
| Assigned To || Due Date || Description || State || Notify || |
Agenda and Minutes
NOTE: Discussion from the telecon is shown in italic.
Calibration Examples Status
Still waiting on final versions of bandpass (RobertLucas
), delay (RobertLucas
), polarization (EdFomalont
), optics (BaltasarVilaVilaro
), and ACA calibration issues (EdFomalont
) calibration examples. See CalExamples
wiki for further information.
Calibration Sequence Development
Will begin process of turning CalExamples
into calibration "sequences" during 2007Q1. Need to first meet with Computing IPT to figure out which Computing IPT personnel will be available to participate in sequence development.
Amplitude Calibration Subsystem Development Reports
- Berkeley Absolute Flux Calibration System (Jack Welch)
- Semi-Transparent Vane (STV) Calibration System
- Need only to get final report from Jesus Martin-Pintado. This is still pending.
- Multi/Dual-Load (M/DL) Calibration System (Matt Carter)
Weather (Ancilliary) Measurement Devices
In December 2006 JeffMangum
submitted weather station instrumentation (RH, P, T, and wind) RFQ documentation to NRAO procurement system. Procurement back in JGM's hands now. After having reviewed information received from bidders, JeffMangum
wonders if it would be more effective to purchase slightly less capable (P,T,RH) but cheaper measurement systems, and to buy more of them to better sample the (P,T,RH) conditions over the ALMA site. Consensus was that this was a good idea. JeffMangum
has posted information on recommended instrumentation to CalAncillary
Based on group discussion JeffMangum will proceed with purchase of proposed (P,T,RH,Ws,Wd) measurement systems.
Allocation of Weather Devices for the OSF
notes that the Science IPT needs to provide a plan, with AIV, for weather station installation at the OSF within the next month or two. At the AOS near the TB in the ESO trailer is a Davis weather station. There is another weather station on a tower near the APEX site, according to Nyman. Neither is currently being used. These could be moved to the holography tower at the OSF.
There are two CR4 hygrometers in the US. One has been repaired by the manufacturer (Buck Research) and was shipped back to Chile last week. Another, at the ATF, is nonfunctional. The repaired instrument will be installed at the OSF.
The CR4 arrived in Santiago and was sent to the OSF. Preliminary discussions were held on its installation on the holography tower. Discussions at the AOS on 2007 Mar 11 confirmed the station on the tower near APEX should be removed, tested and installed near the OSF holography tower. Stations near the containers will remain in place for the time being. Wootten, Nyman, Ocampo and Peck visited the NRAO container and found it to be in reasonable shape. Repair of the 225 GHz tipper was discussed with Donoso. In a short while it could be installed at the AOS and monitored over the internet--this includes leaving it where it is and incorporating it into the AOS infrastructure under operations control.
Solar Observations and Calibration
has written a draft document on Calibration of Solar Observations
. We discussed this document at our last telecon. Few revisions are needed. MarkHoldaway
will update and submit final version within the coming few weeks.
Date of Next Phone Meeting
The next ALMA Calibration Group telecon will be 2007, April 12th 15:00 UT
- 08 Mar 2007